# Quantitative Evaluation of Multiple Instance Learning Reliability For WSIs Classification

### v2

@@ Abstract
[…]
keywords:
Machine learning, Reliability, Whole Slide Image, Multiple Instance Learning, MEAN-POOL-INS.
1 Introduction
@@
Multiple Instance Learning (MIL) is a widely used method for weakly-supervised classification tasks in computational pathology, primarily due to the difficulty and labor-intensive nature of obtaining pixel-wise annotations from histopathology data. In the MIL framework, groups of instances, or "bags," are created, with the goal of predicting the label for each bag. The key assumption in MIL is that positive bags contain at least one key instance, while negative bags do not. Recent advancements in deep learning have spurred the development of novel MIL models tailored to computational pathology, enhancing performance through key instance detection [Ilse et al.(2018)Ilse, Tomczak, and Welling], feature space refinement [Lu et al.(2021)Lu, Williamson, Chen, Chen, Barbieri, and Mahmood], and overfitting mitigation [Zhang et al.(2022)Zhang, Meng, Zhao, Qiao, Yang, Coupland, and Zheng], [Zhang et al.(2023)Zhang, Li, Sun, Zheng, Zhu, and Yang].
@@
Reliability refers to the ability of a system or model to perform as expected under standard and well-defined conditions. The reliability of MIL models is essential for gaining the trust of physicians, ensuring successful deployment in clinical settings, and adhering to established healthcare standards. Reliable models focus on biologically pertinent features, grounded in established scientific knowledge, when making predictions. For instance, in the context of computational pathology, an unreliable MIL model might erroneously focus on artifacts or irrelevant features, leading to false positives or negatives in critical tasks such as tumor detection or subtyping. Such errors could result in misdiagnoses, delays in treatment, or unnecessary interventions, underscoring the need for robust reliability evaluations. Achieving reliability in MIL models is particularly challenging due to the inherent variability in histopathology data, including differences in staining, scanning resolutions, and patient-specific biological heterogeneity. Furthermore, the weakly-supervised nature of MIL exacerbates these challenges by relying on coarse slide-level labels without precise spatial annotations. Addressing these limitations requires not only model improvements but also careful evaluation frameworks that ensure predictions align with domain knowledge and clinical relevance. By fostering reliable MIL models, computational pathology can advance toward broader acceptance and integration into clinical workflows.
@@
However, machine learning models are rarely evaluated in terms of reliability; instead, performance metrics such as accuracy and F1 score are typically reported. Reliable models are easier to deploy in real-world applications and tend to be more generalizable. Therefore, gaining insights into models’ reliability is critical. For instance, Pendyala et al. [Pendyala and Kim(2024)] assessed the reliability of various models using a mental health dataset and discovered that these models often focused on less relevant features, despite achieving high performance, thereby highlighting the limitations of traditional metrics. Furthermore, Araujo et al. [Araujo Santos et al.(2023)Araujo Santos, Cardoso, and Alves] propose an alternative validation method for models and support the assertion that performance metrics are not indicative of the reliability of machine learning networks. In the domain of medical imaging, the trustworthiness of saliency maps has been quantitatively assessed on two public radiology datasets using the area under the precision-recall curve (AUPRC) and the structural similarity index measure (SSIM) as evaluation metrics [Arun et al.(2021)Arun, Gaw, Singh, Chang, Aggarwal, Chen, Hoebel, Gupta, Patel, Gidwani, et al.]. The study revealed that none of the saliency maps satisfied all the tested criteria, highlighting limitations in their reliability.
@@
Interpretability and reliability in the context of machine learning (ML) are related but distinct concepts that address different aspects of model performance and trustworthiness. We aim to understand and explain the way models work. To address questions concerning the decision-making processes and the underlying mechanics, numerous interpretability methods have been developed [Salahuddin et al.(2022)Salahuddin, Woodruff, Chatterjee, and Lambin]. Interpretability and explainability methods have proven useful in identifying both data and model deficiencies [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash]. For MIL models, attention scores assigned to each patch are frequently used to generate heatmaps, which are key tools for assessing interpretability. However, interpretability in MIL models has traditionally been evaluated qualitatively by presenting specific slides and their corresponding heatmaps. The Camelyon16 dataset [Bejnordi et al.(2017)Bejnordi, Veta, Van Diest, Van Ginneken, Karssemeijer, Litjens, Van Der Laak, Hermsen, Manson, Balkenhol, et al.], which includes pixel-wise ground truth annotations, is the most common resource for such evaluations. While qualitative assessments provide valuable insights, they are not feasible across large test sets due to computational and time constraints. Moreover, analyzing heatmaps requires specialized histopathology knowledge, which may not be readily available to machine learning researchers. Finally, a more robust and comprehensive assessment should involve evaluation across multiple datasets.
@@
Reliability is concerned with ensuring the model consistently produces correct and trustworthy predictions in a wide range of circumstances, independent of transparency of underlying process by which predictions are made. Models focusing on related regions of interest (ROIs) are more reliable, as they process the most relevant parts of the data. However, more interpretable models are not necessarily more reliable. For instance, a model might offer more interpretability through class-specific information or excitatory-inhibitory heatmaps [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash], yet fail to concentrate on relevant ROIs. Additionally, while interpretable models may provide explanations for predictions across entire classes, this does not necessarily align with reliability. For example, in binary tumor versus normal classification, since normal samples lack clear ROIs, predictions for these cases are not considered in reliability assessments.
@@
In this paper, we address the above-mentioned challenges by quantitatively evaluating the reliability of MIL models for both binary tumor versus normal classification and multi-class tissue subtyping problems using three metrics. Additionally, we integrate the recently proposed additive method [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash], designed to enhance MIL model interpretability, and evaluate if this also impacts reliability. To provide a more comprehensive assessment, we extend our experiments beyond the Camelyon16 dataset by including two additional datasets: CATCH [Wilm et al.(2022)Wilm, Fragoso, Marzahl, Qiu, Puget, Diehl, Bertram, Klopfleisch, Maier, Breininger, et al.] and TCGA BRCA. Our study aims to advance the development of more reliable MIL models for computational pathology.
2 Method
2.1 Problem formulation
@@
For our analysis, we refer to the input Whole Slide Image (WSI) as $I$ and predict its slide-level label $Y$. Following established methodologies, we first tessellate the slide into $N$ small patches and then extract features from these patches using a pre-trained model. In the final step, the slide label is predicted using an aggregation module, which is typically attention-based, providing attention scores for each patch. For clarification, the proposed framework is illustrated in Figure LABEL:fig:anno.
2.2 Metrics
We propose using three metrics to quantitatively evaluate the reliability of MIL models for WSI classification. Specifically, since our aim is to verify whether MIL models focus on ROIs, we evaluate the alignment between patch scores and annotations from three distinct perspectives:
  • Mutual Information (MI): MI is a measure of statistical dependence. It has been widely used for feature selection and transformation in machine learning [Vergara and Estévez(2014)]. MI quantifies how well values in image $X$ at a given location can predict corresponding values in $Y$. For our purpose, we use MI to evaluate the correspondence of patch scores with their corresponding ground truth.
  • Spearman’s Correlation $(r_{s})$: Spearman’s correlation measures the strength and direction of a non-linear association between two variables. We apply this metric to assess the non-linear correlation between predicted patch scores and ground truth annotations.
  • Area Under The Precision-Recall Curve $(AUPRC)$: AUPRC is particularly useful for evaluating models on highly imbalanced datasets. Given that often our patch annotations are imbalanced, we use AUPRC to measure the localization performance effectively.
3 Experiment and Results
3.1 Dataset
@@
Unlike CAMELYON16 and CATCH, the TCGA BRCA dataset lacks ground truth region-level annotations. To address this, tumor regions in the test slides were manually annotated with the assistance of Dr. Christof A. Bertram, a pathologist.
3.2 Implementation details
Throughout the experiments, we employ the Adam optimizer with a learning rate of 1e-4 and set the maximum number of epochs to 50. The weight decay is tuned based on the validation loss. All experiments are repeated with five different random seeds, and we report the mean AUC and F1 scores as classification metrics, along with the mean MI, $r_{s}$, and AUPRC values as reliability metrics. Additionally, we evaluate model size and Floating-Point Operations Per Second (FLOPS) as computational metrics. For dataset preprocessing, we follow the procedures outlined in [Lu et al.(2021)Lu, Williamson, Chen, Chen, Barbieri, and Mahmood], which involve extracting features from non-overlapping 256×256 patches at 20× magnification using a ResNet50 pretrained on ImageNet. These extracted features are fed into the model with an input dimensionality of $input\_dim=1024$, which is subsequently reduced to $D=512$.
We evaluate several state-of-the-art models, including ABMIL [Ilse et al.(2018)Ilse, Tomczak, and Welling], CLAM [Lu et al.(2021)Lu, Williamson, Chen, Chen, Barbieri, and Mahmood], DTFD [Zhang et al.(2022)Zhang, Meng, Zhao, Qiao, Yang, Coupland, and Zheng], ACMIL [Zhang et al.(2023)Zhang, Li, Sun, Zheng, Zhu, and Yang], MADMIL [Keshvarikhojasteh et al.(2024)Keshvarikhojasteh, Pluim, and Veta], mean pooling (MEAN-POOL), max pooling (MAX-POOL), max pooling instance (MAX-POOL-INS), and mean pooling instance (MEAN-POOL-INS). For the MEAN-POOL and MAX-POOL models, we calculate the slide representation after feature compression using either the mean or max operation, and then predict the slide label. In contrast, for the MAX-POOL-INS and MEAN-POOL-INS models, we first predict patch probabilities, which are then aggregated to generate slide-level probabilities. Patch probabilities are also used to compute reliability metrics for MEAN-POOL-INS. In the case of MAX-POOL-INS, we use both the selected patch index, similar to MAX-POOL, and the patch probabilities.
For multi-head models such as MADMIL and ACMIL, average attention scores or patch probabilities are employed to calculate reliability metrics. We use a shorthand notation to indicate the number of heads, e.g., MADMIL/4 refers to the MADMIL model with four heads. To ensure a thorough analysis, we also investigate additive versions of these models, in accordance with the approach proposed by [Javed et al.(2022)Javed, Juyal, Padigela, Taylor-Weiner, Yu, and Prakash] , which reformulates the final predictor as an additive function of individual instances. For additive models, either attention scores or patch probabilities are used to assess reliability.
@@ 3.2 Implementation details
[…]
fig:anno Refer to caption
Figure 1: The overall framework for evaluating the reliability of MIL models follows a three-step process. First, a MIL model is trained on a weakly-supervised task for predicting slide-level labels. Next, the trained model is applied to predict scores for individual image patches. Finally, the reliability value is computed based on the predicted patch scores and their corresponding annotations, where tumor patches are highlighted in green and normal patches in orange in the annotation visualization.
Table 1: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CAMELYON16 dataset. FLOPs are calculated using 120 instances per bag.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.1166 ± 0.0048 0.2882 ± 0.0044 0.5850 ± 0.0101 0.847 ± 0.011 0.844 ± 0.007 94.4 M 788.7 K CLAM 0.1176 ± 0.0089 0.2925 ± 0.0173 0.5889 ± 0.0220 0.860 ± 0.005 0.850 ± 0.007 94.4 M 790.7 K DTFD 0.1275 ± 0.0071 0.2939 ± 0.0106 0.6057 ± 0.0142 0.893 ± 0.011 0.847 ± 0.017 125.8 M 1052.6 K ACMIL/2 0.1612 ± 0.0198 0.3352 ± 0.0168 0.5856 ± 0.0486 0.824 ± 0.028 0.810 ± 0.014 94.4 M 791.0 K MADMIL/2 0.1330 ± 0.0045 0.3057 ± 0.0043 0.6080 ± 0.0037 0.862 ± 0.009 0.847 ± 0.013 78.6 M 657.6 K MAX-POOL 0.0018 ± 0.0001 -$0$ 0.0014 0.2253 ± 0.0097 0.798 ± 0.006 0.793 ± 0.012 62.9 M 525.8 K MEAN-POOL [^1] The reliability metrics are not computed for this model, as it does not provide patch scores. - - - 0.581 ± 0.022 0.589 ± 0.021 62.9 M 525.8 K
[…]
fig:cam_max Refer to caption        Refer to caption     Refer to caption
Figure 2: The test-30 slide with ground truth annotation (green) along with corresponding heatmap generated by MAX-POOL.
Table 2: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CAMELYON16 using additive models.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.1336 ± 0.0031 0.3179 ± 0.0081 0.6152 ± 0.0107 0.861 ± 0.010 0.827 ± 0.015 94.5 M 788.7 K ABMIL-ADD-PATCH 0.0380 ± 0.0025 0.2377 ± 0.0112 0.4848 ± 0.0070 CLAM-ADD-ATT 0.1153 ± 0.0091 0.2924 ± 0.0089 0.5968 ± 0.0170 0.865 ± 0.006 0.826 ± 0.019 94.5 M 790.7 K CLAM-ADD-PATCH 0.0325 ± 0.0042 0.2394 ± 0.0499 0.4826 ± 0.0165 DTFD-PATCH  [^2] We use the provided patch probabilities based on GRAD-CAM. 0.0368 ± 0.0030 0.2895 ± 0.0197 0.5308 ± 0.0124 0.893 ± 0.011 0.847 ± 0.017 125.8 M 1052.6 K ACMIL/2-ADD-ATT 0.1739 ± 0.0055 0.3463 ± 0.0117 0.6201 ± 0.0220 0.787 ± 0.024 0.810 ± 0.011 94.5 M 791.0 K ACMIL/2-ADD-PATCH 0.0263 ± 0.0033 -0.0346 ± 0.0120 0.2873 ± 0.0088 MAX-POOL-INS 0.0004 ± 0.0002 0.0473 ± 0.0255 0.1986 ± 0.0228 0.766 ± 0.135 0.763 ± 0.127 63.0 M 525.8 K MAX-POOL-INS-PATCH 0.1498 ± 0.0561 0.2932 ± 0.1501 0.5536 ± 0.1972 MEAN-POOL-INS 0.3097 ± 0.0367 0.5338 ± 0.0678 0.7841 ± 0.0780 0.581 ± 0.022 0.589 ± 0.021 63.0 M 525.8 K
3.3 Results
3.3.1 CAMELYON16

### v3

# Beyond accuracy: quantifying the reliability of Multiple Instance Learning for Whole Slide Image classification

@@ Abstract
Machine learning models have become integral to many fields, but their reliability, defined as producing dependable, trustworthy, and domain-consistent predictions, remains a critical concern. Multiple Instance Learning (MIL) models designed for Whole Slide Image (WSI) classification in computational pathology are rarely evaluated in terms of reliability, leaving a key gap in understanding their suitability for high-stakes applications like clinical decision-making. In this paper, we address this gap by introducing three quantitative metrics for reliability assessment and applying them to several widely used MIL architectures across three region-wise annotated pathology datasets. Our findings indicate that the mean pooling instance (MEAN-POOL-INS) model demonstrates superior reliability compared to other networks, despite its simple architectural design and computational efficiency. These findings underscore the need of reliability evaluation alongside predictive performance in MIL models and establish MEAN-POOL-INS as a strong, trustworthy baseline for future research.
Introduction
Machine learning (ML) has become a cornerstone of modern computational pathology by enabling automated analysis of large and complex datasets generated in clinical practice [1, 2]. Deep learning, in particular, has shown remarkable potential for tasks such as tumor detection, tissue subtyping, and prognosis prediction [3, 4].
@@ Introduction
Multiple Instance Learning (MIL) is a widely used method for weakly-supervised classification tasks in computational pathology, primarily due to the difficulty and labor-intensive nature of obtaining pixel-wise annotations from histopathology data [5, 6]. Compared to simpler approaches such as patch-level classifiers with propagated slide labels or heuristic aggregation strategies, MIL offers a principled means of reducing label noise and focusing on diagnostically relevant regions. In the MIL framework, groups of instances, or ”bags,” are created, with the goal of predicting the label for each bag. The key assumption in MIL is that positive bags contain at least one key instance, while negative bags do not. Recent advancements in deep learning have spurred the development of novel MIL models tailored to computational pathology, enhancing performance through key instance detection [7], feature space refinement [8], and overfitting mitigation [9, 10].
@@
Reliability refers to the ability of a system or model to perform as expected under standard and well-defined conditions. The reliability of MIL models is essential for gaining the trust of physicians, ensuring successful deployment in clinical settings, and adhering to established healthcare standards. Reliable models focus on biologically pertinent features, grounded in established scientific knowledge, when making predictions. In the context of computational pathology, an unreliable MIL model might erroneously focus on artifacts or irrelevant features, leading to false positives or negatives in critical tasks such as tumor detection or subtyping. For example, a model could achieve high slide-level accuracy by correctly predicting tumor presence in most cases, yet consistently rely on irrelevant regions such as staining or scanner-induced artifacts, making it accurate but not reliable. Conversely, a model could consistently focus on biologically meaningful regions (reliable) but exhibit slightly lower overall accuracy due to ambiguous or noisy labels. Such errors could result in misdiagnoses, delays in treatment, or unnecessary interventions, underscoring the need for robust reliability evaluations. Achieving reliability in MIL models for computational patholopgy is particularly challenging due to the inherent variability in histopathology data, including differences in staining, scanning resolutions, and patient-specific biological heterogeneity [11]. Furthermore, the weakly-supervised nature of MIL exacerbates these challenges by relying on coarse slide-level labels without precise spatial annotations. Addressing these limitations requires not only model improvements but also careful evaluation frameworks that ensure predictions align with domain knowledge and clinical relevance. By fostering reliable MIL models, computational pathology can advance toward broader acceptance and integration into clinical workflows.
@@
However, machine learning models are rarely evaluated in terms of reliability; instead, performance metrics such as accuracy and F1 score are typically reported [12]. Yet, reliable models are essential for real-world deployment and tend to exhibit better generalization. Therefore, gaining insights into models’ reliability is critical. For instance, Pendyala et al. [13] assessed the reliability of various models using a mental health dataset and discovered that these models often focused on less relevant features, despite achieving high performance, thereby highlighting the limitations of traditional metrics. Furthermore, Araujo et al. [14] propose an alternative validation method for models and support the assertion that performance metrics are not indicative of the reliability of machine learning networks. In the domain of medical imaging, the trustworthiness of saliency maps has been quantitatively assessed on two public radiology datasets using the area under the precision-recall curve (AUPRC) and the structural similarity index measure (SSIM) as evaluation metrics [15]. The study revealed that none of the saliency maps satisfied all the tested criteria, highlighting limitations in their reliability.
@@
Interpretability and reliability in the context of machine learning (ML) are related but distinct concepts that address different aspects of model performance and trustworthiness. Interpretability means how easily a person can understand why a model made a certain decision or how it comes to its results [16]. We aim to understand and explain the way models work in interpretability, providing transparency. To address questions concerning the decision-making processes and the underlying mechanics, numerous interpretability methods have been developed [17]. Interpretability and explainability methods have proven useful in identifying both data and model deficiencies [18]. For MIL models, attention scores assigned to each patch are frequently used to generate heatmaps, which are key tools for assessing interpretability. However, interpretability in MIL models has traditionally been evaluated qualitatively by presenting specific slides and their corresponding heatmaps. The Camelyon16 dataset [19], which includes pixel-wise ground truth annotations, is the most common resource for such evaluations. While qualitative assessments provide valuable insights, they are not feasible across large test sets due to computational and time constraints. Moreover, analyzing heatmaps requires specialized histopathology knowledge, which may not be readily available to machine learning researchers. Finally, a more robust and comprehensive assessment should involve evaluation across multiple datasets.
@@
In contrast to interpretability, reliability is concerned with ensuring the model consistently produces correct and trustworthy predictions in a wide range of circumstances, independent of transparency of underlying process by which predictions are made. Models focusing on related regions of interest (ROIs) are more reliable, as they process the most relevant parts of the data. However, more interpretable models are not necessarily more reliable. For instance, a model might offer more interpretability through class-specific information or excitatory-inhibitory heatmaps [18], yet fail to concentrate on relevant ROIs. Additionally, while interpretable models may provide explanations for predictions across entire classes, this does not necessarily align with reliability. For example, in binary tumor versus normal classification, since normal samples lack clear ROIs, predictions for these cases are not considered in reliability assessments.
@@
In this paper, we address the above-mentioned challenges by quantitatively evaluating the reliability of MIL models for both binary tumor versus normal classification and multi-class tissue subtyping problems using three metrics. Additionally, we incorporate a recently proposed additive interpretability method [18] to investigate whether improved interpretability translates into increased reliability, providing new insights into the complex relationship between interpretability and reliability. Unlike previous studies that rely mainly on single-dataset evaluations, we extend our experiments beyond the Camelyon16 dataset by including two additional datasets: CATCH [20] and TCGA BRCA. By introducing quantitative reliability metrics, benchmarking multiple MIL architectures across three datasets, and examining the impact of additive interpretability, this study provides a comprehensive, multi-dataset evaluation of MIL model reliability in computational pathology. Our study aims to advance the development of more reliable MIL models for computational pathology.
The main contributions of this paper are as follows:
  • Introduction of three quantitative metrics to evaluate the reliability of MIL models in computational pathology.
  • Systematic benchmarking of several widely-used MIL architectures across three region-wise annotated pathology datasets: Camelyon16, CATCH, and TCGA BRCA.
  • Investigation of the impact of a recently proposed additive interpretability method on MIL model reliability, providing insights into the relationship between interpretability and reliability.
  • Demonstration that comprehensive reliability evaluation complements traditional accuracy metrics and is essential for trustworthy deployment in high-stakes clinical settings.
Materials and methods
In this section, we describe the problem formulation, the evaluation metrics for predictive performance and reliability, the datasets used in our experiments, and the details of our implementation.
Problem formulation
@@
For our analysis, we refer to the input Whole Slide Image (WSI) as $I$ and predict its slide-level label $\hat{Y}$. Following established methodologies, we first tessellate the slide into $N$ small patches and then extract features from these patches using a pre-trained model. In the final step, the slide label is predicted using an aggregation module, which is typically attention-based, providing attention scores for each patch. For clarification, the proposed framework is illustrated in Fig 1.
Fig 1: The overall framework for evaluating the reliability of MIL models follows a three-step process. First, a MIL model is trained on a weakly-supervised task for predicting slide-level labels. Next, the trained model is applied to predict scores for individual image patches. Finally, the reliability value is computed based on the predicted patch scores and their corresponding annotations, where tumor patches are highlighted in green and normal patches in orange in the annotation visualization.
Metrics
Reliability in this study is defined as the consistent focus of MIL models on diagnostically relevant ROIs within WSIs, a prerequisite for trustworthy and clinically useful predictions. To quantify this alignment, we selected three complementary metrics that together capture different aspects of spatial concordance between predicted patch scores and ground truth annotations:
  • Mutual Information $(MI)$: Measures the statistical dependence between predicted scores and ROIs, quantifying how much information the model’s output shares with the true ROI distribution [21].
  • Spearman’s Correlation $(Spearman\text{’}s)$: Captures the monotonic relationship between predicted patch importance and true ROI presence, allowing for non-linear associations and reflecting relative ranking fidelity [22].
  • Area Under the Precision-Recall Curve $(AUPRC)$: Evaluates the model’s ability to localize ROIs under conditions of class imbalance by measuring precision and recall trade-offs [23].
Dataset
@@
Unlike CAMELYON16 and CATCH, the TCGA BRCA dataset lacks ground truth region-level annotations. To address this, tumor regions in the test slides were manually annotated with the assistance of Dr. Christof A. Bertram, a pathologist [24].
Implementation details
Throughout the experiments, we employ the Adam optimizer with a learning rate of $1\times 10^{-4}$ and set the maximum number of epochs to 50. The weight decay is tuned based on the validation loss. All experiments are repeated with five different random seeds, and we report the mean AUC and F1 scores as classification metrics, along with the mean MI, $r_{s}$, and AUPRC values as reliability metrics. The AUC measures the model’s overall discriminative ability in classifying positive and negative cases across all thresholds, while the AUPRC emphasizes the model’s reliability in identifying true positive regions under conditions of class imbalance. Using AUPRC as a reliability metric thus complements the classification-based AUC, providing a more comprehensive evaluation of MIL model performance. Additionally, we evaluate model size and Floating-Point Operations Per Second (FLOPS) as computational metrics. For dataset preprocessing, we follow the procedures outlined in [8], which involve extracting features from non-overlapping $256\times 256$ patches at $20\times$ magnification using a ResNet50 pretrained on ImageNet. These extracted features are fed into the model with an input dimensionality of input_dim=1024input\_dim=1024, which is subsequently reduced to $D=512$.
We evaluate several state-of-the-art models, including ABMIL [7], CLAM [8], DTFD [9], ACMIL [10], MADMIL [25], mean pooling (MEAN-POOL), max pooling (MAX-POOL), max pooling instance (MAX-POOL-INS), and mean pooling instance (MEAN-POOL-INS). For the MEAN-POOL and MAX-POOL models, we calculate the slide representation after feature compression using either the mean or max operation, and then predict the slide label. It should be noted that reliability metrics are not reported for the MEAN-POOL model, as its mean aggregation does not yield patch-level scores or attention weights. In contrast, for the MAX-POOL-INS and MEAN-POOL-INS models, we first predict patch probabilities, which are then aggregated to generate slide-level probabilities. Patch probabilities are also used to compute reliability metrics for MEAN-POOL-INS. In the case of MAX-POOL-INS, we use both the selected patch index, similar to MAX-POOL, and the patch probabilities.
For multi-head models such as MADMIL and ACMIL, average attention scores or patch probabilities are employed to calculate reliability metrics. We use a shorthand notation to indicate the number of heads, e.g., MADMIL/4 refers to the MADMIL model with four heads. To ensure a thorough analysis, we also investigate additive versions of these models, in accordance with the approach proposed by [18] , which reformulates the final predictor as an additive function of individual instances. For additive models, either attention scores or patch probabilities are used to assess reliability.
@@
Results
We present the predictive performance and reliability results for CAMELYON16, CATCH, and TCGA BRCA datasets. This section highlights how well each model identifies regions of interest and compares their overall effectiveness.
CAMELYON16